• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name E.Chudakov

    Log entry time 00:23:45 on November30,2006

    Entry number 189514

    keyword=Shift summary

    No beam till about 18:00 (spin dance work)
    17:10 - control access to investigate a possible target leak (JP)
    Beam came at about 20:00
    At 21:30 an accelerator vacuum problem occurred, then an optics problem did not allow to send us a beam.

    1139 - elastic, starting values
    1140 - elastic, HV pion rejector changed (/adaqfs/home/adev/dv_rp.3
    1141 - inelastic (HRS L moved to 1.06GeV/c)
    1143 - New voltage on gas Cherenkov